OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. D(5, -2) m = 2/5

OpenStudy (anonymous):

Start with the point-slope form of a line, then convert to slope intercept. \[y-y_0 = m(x-x_0)\] Where \(x_0,y_0\) is the x,y values of the point and \(m\) is the slope. Then convert to \(y=mx+b\) form. \[y-y_0 = m(x-x_0)\] \[\implies y = mx +(y_0-mx_0)\]

OpenStudy (anonymous):

wait so would it be y=2/5x+(-2-(2/5)*5))?

OpenStudy (anonymous):

After you simplify, yes.

OpenStudy (anonymous):

\[y=\frac{2}{5}x + (-2-(2/5)*5) \] \[=\frac{2}{5}x +(-2-2)\] \[=\frac{2}{5}x -4\]